# Discrete Guidance Matching: Exact Guidance for Discrete Flow Matching

### v1

@@ Appendix B Additional Related Works
Discrete Flow-based Models. Flow-based models for modeling discrete data were introduced by [9, 22, 65, 28, 58]. Compared to discrete diffusion models [3, 8, 71, 47], the main advantage of discrete flow models is the flexibility of designing the conditional probability path and the conditional transition rate without specifying corruption transition rate; while discrete diffusion models require choosing a forward transition rate with a simple form such that the conditional probability path can be computed easily. Similar to the framework of continuous flow models [1, 45, 40], discrete flow models focus on the conditional path with probability interpolation, which is a convex combination of conditional probabilities. [65] proposed the mixture probability path with $x_{1}$-dependent schedulers and the metric-induced path; the training objective they used is the negative ELBO instead of the cross entropy used in [9] and [22]. Recently, [25] introduced a novel discrete flow matching framework with auxiliary process, which supports variable-length sequence generation.
@@
Discrete Diffusion Models. Motivated by studies of discrete-time diffusion models in continuous state spaces [68, 26, 69, 33], [3] and [30] proposed a discrete-time framework for training diffusion models in discrete state spaces. Similarly, several authors have developed continuous-time discrete diffusion models through continuous-time Markov chains [8, 6]. [71] proposed discrete score matching based on the idea of continuous score matching for modeling continuous-time diffusion models [70]. [49, 47] introduced concrete score matching, which aims to identify the time-reversal of the noising process like [71]. Empirically, the masked diffusion models [66, 63, 54] outperform the diffusion models with other noise distributions, such as uniform distribution. Moreover, for masked diffusion models, [54] showed that the training objective is equivalent to that of any-order autoregressive models [29, 67], which is also equivalent to the ELBO for discrete flow [see Appendix D.1 in 65]. The main advantage of discrete diffusion models over (any-order) autoregressive models is parallel sampling. However, supporting variable-length sequence generation is considerably more challenging than in autoregressive models. To overcome this limitation, semi-autoregressive models [51, 2] perform block-wise sampling and enable parallel sampling within each block, thus supporting flexible-length sampling.
@@
Guidance. Classifier guidance [16] and classifier-free guidance [27, 88] are proposed for conditional generation. [64] explored the classifier guidance tailored to discrete diffusion models by decomposing the reversed transition probability for practical purposes. [50] applied classifier-free guidance to masked diffusion models for unsupervised pretraining. In this paper, the posterior-based guidance provides a unified perspective for the existing guidance in previous work, including classifier guidance [16], energy-weighted guidance [48], guidance for transfer learning [56], and flow matching guidance [20] in continuous state space. See discussion in Section E.1. Moreover, thanks to the token-wise transition rate, our discrete guidance only requires training one network, unlike the training-based continuous guidance proposed by [20], which requires estimating a normalizing constant by learning a surrogate model.
@@
Diffusion-based and Flow-based Large Language Models. Recently, a number of studies have focused on developing large language diffusion models. Several post-training approaches have been proposed to fine-tune discrete diffusion models aligned with human preference in a way similar to autoregressive models via RLHF [55] or DPO [60] techniques. LLaDa [51] leverages masked diffusion models and employs negative ELBO for supervised fine-tuning. It supports variable-length sentence generation by semi-autoregressive, and utilizes a deterministic low-confidence remasking technique to enhance performance similar to MaskGIT [10]. LLaDa 1.5 [90] applied DPO to LLaDa, using an empirical estimator with variance reduction techniques to approximate the log-probability. To extend GRPO to masked diffusion models, d1-LLaDa [87] proposes to estimate log-probability via mean-field approximation and to do one-step unmasking to estimate token-wise log-probability with partially masked prompts. To learn from multi-step denoising information, MMaDa [83] perturbs sentence outcomes instead of prompts, and conducts policy optimization with a GRPO-type objective, which supports multi-modality. For the large language model based on discrete flow matching, [78] developed a multi-modal model with a metric-induced probability path and a kinetic-optimal conditional rate [65], achieving highly competitive performance compared to advanced AR-based multi-modal models.
@@ Appendix C Background of Continuous-Time Markov Chains
Intuitively, the transition rate $u_{t}(z,x)$ measures the intensity of transition from the current state $x$ to the target state $z$ at time $t$. This CTMC framework underlies recent advances in generative modeling. Building on the framework of CTMCs, the discrete diffusion models [3, 8, 71, 47] and the discrete flow models [9, 22, 65, 58] can transport a simple distribution like uniform distribution or just a point mass to a complex data distribution through a sequence of jumps.
@@ Appendix D Derivation of the guidance
D.1 Proof of Proposition 3.1 of [9]
We include the proof of Proposition 3.1 of [9] here for completion. We begin by taking the expectation over $p_{1}$ on both sides of the Kolmogorov equation for $p_{t|1}(x|x_{1})$ and $u_{t}(x_{t},z|x_{1})$.
$\partial_{t}p_{t}(x)$ $=\mathbb{E}_{{\mathbf{x}}\sim p_{1}({\mathbf{x}}_{1})}\big[\partial_{t}p_{t|1}(x|{\mathbf{x}}_{1})\big]\qquad{\small\text{(definition of $p_{t}$ as marginal over $x_{1}$)}}$
$=\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1}(x_{1})}\big[\sum_{z}u_{t}(x,z|{\mathbf{x}}_{1})\,p_{t|1}(z|{\mathbf{x}}_{1})\big]\qquad{\small\text{(Kolmogorov forward equation)}}$
@@
$=\sum_{z}\sum_{x_{1}}p_{t}(z)\,p_{1|t}(x_{1}|z)\,u_{t}(x,z|x_{1})\qquad{\small\text{(Bayes' rule)}}$
@@ E.1.1 Continuous Score Matching
where we use the facts $q_{t|0}=p_{t|0}$ and $\mathbb{E}_{{\mathbf{x}}_{0}\sim p_{0|t}({\mathbf{x}}_{0}|x)}{[r({\mathbf{x}}_{0})]}=\frac{q_{t}(x)}{p_{t}(x)}$. This result recovers the guidance given by [56]. By setting $q_{t}(x)=p_{t}(x|y)$, we also recover the classifier guidance $\nabla_{x}\log q_{t}(x)=\nabla_{x}\log p_{t}(x)+\nabla_{x}\log p(y|{\mathbf{x}}_{t}=x)$.
@@ E.1.2 Continuous Flow Matching
Denote $v_{t}(x)$ and $v_{t}(x|x_{1})$ as marginal velocity and conditional velocity, respectively. Note that
$v_{t}^{q}(x)=$ $~\mathbb{E}_{{\mathbf{x}}_{1}\sim q_{1|t}({\mathbf{x}}_{1}|x)}{[v_{t}(x|{\mathbf{x}}_{1})]}$
$=$ $~\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|x)}{[v_{t}(x|{\mathbf{x}}_{1})\frac{q_{1|t}({\mathbf{x}}_{1}|x)}{p_{1|t}({\mathbf{x}}_{1}|x)}]}$
$=$ $~v_{t}^{p}(x)+\int_{x_{1}}{\Big(\frac{r(x_{1})}{\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|x)}{[r({\mathbf{x}}_{1})]}}-1\Big)}v_{t}(x|x_{1})p_{1|t}(x_{1}|x)\text{d}x_{1},$
which coincides the exact guidance [20] for continuous flow matching.
@@ E.1.4 Discrete Flow Matching
To give a specific example, we consider the $x_{1}$-independent mixture path and its kinetic-optimal transition rate [see Appendix C of 65]:
@@ E.2 Discrete Guidance for Discrete-Time Framework
In this subsection, we discuss the discrete guidance in the settings of the discrete-time Markov chain [e.g., D3PM, 3], which is developed by [76, 64, 37]. For convenience, we also use $t=1$ for data distribution and $t=0$ for initial distribution. Assume that $p_{s|t}=q_{s|t}$ for each $s<t$. By Equation 16, the transition probability of $q$ in the sampling process is
@@
which is similar to the soft optimal policy introduced in [37]. By taking $q_{t+h|t}(z|x)=p_{t+h|t}(z|x,y)$, we recover the classifier guidance in [76] and [64] with unit guidance strength.
Following the proof of Theorem 2, the rate-based guidance can be viewed as a limiting form of the discrete guidance in discrete time settings. In addition, the discrete-time guidance has a similar challenge to the rate-based guidance; that is, it requires multiple function calls in each sampling step. For solving this problem, [76, 64] employ the first-order approximation in practice. As pointed out in Remark 1, such approximation-based approaches are inappropriate for discrete data intuitively.
@@ E.3 Comparison between Posterior-Based Guidance and Rate-Based Guidance
Theorem 1 and Theorem 2 present the discrete guidance based on the posterior and the transition rate, respectively. The later one generalizes the discrete guidance of [52], which requires a stronger condition than the posterior-based guidance in Theorem 1. To better understand the conditions in both theorems, we will analyze the relationship between discrete diffusion and discrete flow models. In the following, we consider two different states $x\neq z$.
Both Guidance Yield the Same Transition Rates in Discrete Diffusion Models. Consider a discrete diffusion model with the corruption transition rate $Q^{q}_{t}(x,z)$. Here, we assume that $q_{1}$ is the density of the data distribution. Following [8], the time reversal of $Q^{q}_{t}(x,z)$ is
@@
Since $u_{t}^{q}(z,x|x_{1})\overset{\triangle}{=}Q^{q}_{t}(x,z)\frac{q_{t|1}(z|x_{1})}{q_{t|1}(x|x_{1})}$ is a conditional transition rate that can generate the conditional path $q_{t|1}$ [see Appendix H.1 of 9], the discrete diffusion models can be viewed as a special case of discrete flow models. Therefore, the condition $Q^{q}_{t}(x,z)=Q^{p}_{t}(x,z)$ as in Theorem 2 actually implies that we take the same specific conditional transition rate $u_{t}^{p}(z,x|x_{1})=u_{t}^{q}(z,x|x_{1})=Q^{q}_{t}(x,z)\frac{q_{t|1}(z|x_{1})}{q_{t|1}(x|x_{1})}$ in discrete flow models.
They Might Offer Different Marginal Transition Rates for Sampling from Target Distribution. Suppose that we are given conditional probability path $q_{t|1}=p_{t|1}$ and conditional transition rate $u_{t}^{q}(z,x|x_{1})=u_{t}^{p}(z,x|x_{1})$. What we want to know is whether there exists a common corruption transition rate $Q_{t}(x,z)$ such that $u^{q}_{t}(z,x)$ and $u^{p}_{t}(z,x)$ are the reverse transition rates of $Q_{t}(x,z)$ under the data distributions $q_{1}$ and $p_{1}$, respectively. If exists, by the time reversal formula [see, e.g. 71, 47], we have
@@
where the third equation we use $\frac{q_{1|t}({\mathbf{x}}_{1}|x)}{p_{1|t}({\mathbf{x}}_{1}|x)}=\frac{q_{1}({\mathbf{x}}_{1})p_{t}(x)}{p_{1}({\mathbf{x}}_{1})q_{t}(x)}$ in equation 15 and the fourth equation we use the marginalization trick in the concrete score matching [49, 47]. Then we are done.

### v2

@@ Appendix B Additional Related Works
Discrete Flow-based Models. Flow-based models for modeling discrete data were introduced by Campbell et al. (2024); Gat et al. (2024); Shaul et al. (2025); Holderrieth et al. (2025); Qin et al. (2025). Compared to discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Sun et al., 2023; Lou et al., 2024), the main advantage of discrete flow models is the flexibility of designing the conditional probability path and the conditional transition rate without specifying corruption transition rate; while discrete diffusion models require choosing a forward transition rate with a simple form such that the conditional probability path can be computed easily. Similar to the framework of continuous flow models (Albergo and Vanden-Eijnden, 2023; Liu et al., 2022; Lipman et al., 2023), discrete flow models focus on the conditional path with probability interpolation, which is a convex combination of conditional probabilities. Shaul et al. (2025) proposed the mixture probability path with $x_{1}$-dependent schedulers and the metric-induced path; the training objective they used is the negative ELBO instead of the cross entropy used in Campbell et al. (2024) and Gat et al. (2024). Recently, Havasi et al. (2025) introduced a novel discrete flow matching framework with auxiliary process, which supports variable-length sequence generation.
@@
Discrete Diffusion Models. Motivated by studies of discrete-time diffusion models in continuous state spaces (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021a; Karras et al., 2022), Austin et al. (2021) and Hoogeboom et al. (2021) proposed a discrete-time framework for training diffusion models in discrete state spaces. Similarly, several authors have developed continuous-time discrete diffusion models through continuous-time Markov chains (Campbell et al., 2022; Benton et al., 2024). Sun et al. (2023) proposed discrete score matching based on the idea of continuous score matching for modeling continuous-time diffusion models (Song et al., 2021b). Meng et al. (2022); Lou et al. (2024) introduced concrete score matching, which aims to identify the time-reversal of the noising process like Sun et al. (2023). Empirically, the masked diffusion models (Shi et al., 2024; Sahoo et al., 2024; Ou et al., 2025) outperform the diffusion models with other noise distributions, such as uniform distribution. Moreover, for masked diffusion models, Ou et al. (2025) showed that the training objective is equivalent to that of any-order autoregressive models (Hoogeboom et al., 2022; Shih et al., 2022), which is also equivalent to the ELBO for discrete flow (see Appendix D.1 in Shaul et al., 2025). The main advantage of discrete diffusion models over (any-order) autoregressive models is parallel sampling. However, supporting variable-length sequence generation is considerably more challenging than in autoregressive models. To overcome this limitation, semi-autoregressive models (Nie et al., 2025b; Arriola et al., 2025) perform block-wise sampling and enable parallel sampling within each block, thus supporting flexible-length sampling.
@@
Guidance. Classifier guidance (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2021; Zheng et al., 2023) are proposed for conditional generation. Schiff et al. (2025) explored the classifier guidance tailored to discrete diffusion models by decomposing the reversed transition probability for practical purposes. Nie et al. (2025a) applied classifier-free guidance to masked diffusion models for unsupervised pretraining. In this paper, the posterior-based guidance provides a unified perspective for the existing guidance in previous work, including classifier guidance Dhariwal and Nichol (2021), energy-weighted guidance Lu et al. (2023), guidance for transfer learning (Ouyang et al., 2024), and flow matching guidance (Feng et al., 2025) in continuous state space. See discussion in Section E.1. Moreover, thanks to the token-wise transition rate, our discrete guidance only requires training one network, unlike the training-based continuous guidance proposed by (Feng et al., 2025), which requires estimating a normalizing constant by learning a surrogate model.
@@
Diffusion-based and Flow-based Large Language Models. Recently, a number of studies have focused on developing large language diffusion models. Several post-training approaches have been proposed to fine-tune discrete diffusion models aligned with human preference in a way similar to autoregressive models via RLHF (Ouyang et al., 2022) or DPO (Rafailov et al., 2023) techniques. LLaDa (Nie et al., 2025b) leverages masked diffusion models and employs negative ELBO for supervised fine-tuning. It supports variable-length sentence generation by semi-autoregressive, and utilizes a deterministic low-confidence remasking technique to enhance performance similar to MaskGIT (Chang et al., 2022). LLaDa 1.5 (Zhu et al., 2025) applied DPO to LLaDa, using an empirical estimator with variance reduction techniques to approximate the log-probability. To extend GRPO to masked diffusion models, d1-LLaDa (Zhao et al., 2025) proposes to estimate log-probability via mean-field approximation and to do one-step unmasking to estimate token-wise log-probability with partially masked prompts. To learn from multi-step denoising information, MMaDa Yang et al. (2025) perturbs sentence outcomes instead of prompts, and conducts policy optimization with a GRPO-type objective, which supports multi-modality. For the large language model based on discrete flow matching, Wang et al. (2025) developed a multi-modal model with a metric-induced probability path and a kinetic-optimal conditional rate (Shaul et al., 2025), achieving highly competitive performance compared to advanced AR-based multi-modal models.
@@ Appendix C Background of Continuous-Time Markov Chains
Intuitively, the transition rate $u_{t}(z,x)$ measures the intensity of transition from the current state $x$ to the target state $z$ at time $t$. This CTMC framework underlies recent advances in generative modeling. Building on the framework of CTMCs, the discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Sun et al., 2023; Lou et al., 2024) and the discrete flow models (Campbell et al., 2024; Gat et al., 2024; Shaul et al., 2025; Qin et al., 2025) can transport a simple distribution like uniform distribution or just a point mass to a complex data distribution through a sequence of jumps.
@@ Appendix D Derivation of the guidance
D.1 Proof of Proposition 3.1 of (Campbell et al., 2024)
We include the proof of Proposition 3.1 of (Campbell et al., 2024) here for completion. We begin by taking the expectation over $p_{1}$ on both sides of the Kolmogorov equation for $p_{t|1}(x|x_{1})$ and $u_{t}(x_{t},z|x_{1})$.
$\partial_{t}p_{t}(x)$ $=\mathbb{E}_{{\mathbf{x}}\sim p_{1}({\mathbf{x}}_{1})}\big[\partial_{t}p_{t|1}(x|{\mathbf{x}}_{1})\big]\quad{\text{(definition of $p_{t}$ as marginal over $x_{1}$)}}$
$=\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1}(x_{1})}\big[\sum_{z}u_{t}(x,z|{\mathbf{x}}_{1})\,p_{t|1}(z|{\mathbf{x}}_{1})\big]\quad{\text{(Kolmogorov forward equation)}}$
@@
$=\sum_{z}\sum_{x_{1}}p_{t}(z)\,p_{1|t}(x_{1}|z)\,u_{t}(x,z|x_{1})\quad{\text{(Bayes' rule)}}$
@@ E.1.1 Continuous Score Matching
where we use the facts $q_{t|0}=p_{t|0}$ and $\mathbb{E}_{{\mathbf{x}}_{0}\sim p_{0|t}({\mathbf{x}}_{0}|x)}{[r({\mathbf{x}}_{0})]}=\frac{q_{t}(x)}{p_{t}(x)}$. This result recovers the guidance given by Ouyang et al. (2024). By setting $q_{t}(x)=p_{t}(x|y)$, we also recover the classifier guidance $\nabla_{x}\log q_{t}(x)=\nabla_{x}\log p_{t}(x)+\nabla_{x}\log p(y|{\mathbf{x}}_{t}=x)$.
@@ E.1.2 Continuous Flow Matching
Denote $v_{t}(x)$ and $v_{t}(x\mid x_{1})$ as the marginal and conditional velocity fields, respectively. Assuming the conditional kernels satisfy $q_{t\mid 1}(x\mid x_{1})=p_{t\mid 1}(x\mid x_{1})$, and defining $r(x_{1}):=q_{1}(x_{1})/p_{1}(x_{1})$, we have
$v_{t}^{q}(x)$ $=\mathbb{E}_{{\mathbf{x}}_{1}\sim q_{1\mid t}({\mathbf{x}}_{1}\mid x)}{[v_{t}(x\mid{\mathbf{x}}_{1})]}$
$=\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1\mid t}({\mathbf{x}}_{1}\mid x)}{[v_{t}(x\mid{\mathbf{x}}_{1})\frac{q_{1\mid t}({\mathbf{x}}_{1}\mid x)}{p_{1\mid t}({\mathbf{x}}_{1}\mid x)}]}$
$=\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1\mid t}({\mathbf{x}}_{1}\mid x)}{[v_{t}(x\mid{\mathbf{x}}_{1})\frac{r({\mathbf{x}}_{1})}{\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1\mid t}({\mathbf{x}}_{1}\mid x)}[r({\mathbf{x}}_{1})]}]}$
$=v_{t}^{p}(x)+\int_{x_{1}}{\Big(\frac{r(x_{1})}{\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1\mid t}({\mathbf{x}}_{1}\mid x)}[r({\mathbf{x}}_{1})]}-1\Big)}v_{t}(x\mid x_{1})\,p_{1\mid t}(x_{1}\mid x)\,\text{d}x_{1}.$
which coincides the exact guidance (Feng et al., 2025) for continuous flow matching.
@@ E.1.4 Discrete Flow Matching
To give a specific example, we consider the $x_{1}$-independent mixture path and its kinetic-optimal transition rate (see Appendix C of Shaul et al., 2025):
@@ E.2 Discrete Guidance for Discrete-Time Framework
In this subsection, we discuss the discrete guidance in the settings of the discrete-time Markov chain (e.g., D3PM, Austin et al., 2021), which is developed by Vignac et al. (2023); Schiff et al. (2025); Li et al. (2024a). For convenience, we also use $t=1$ for data distribution and $t=0$ for initial distribution. Assume that $p_{s|t}=q_{s|t}$ for each $s<t$. By Equation 16, the transition probability of $q$ in the sampling process is
@@
which is similar to the soft optimal policy introduced in Li et al. (2024a). By taking $q_{t+h|t}(z|x)=p_{t+h|t}(z|x,y)$, we recover the classifier guidance in Vignac et al. (2023) and Schiff et al. (2025) with unit guidance strength.
Following the proof of Theorem 2, the rate-based guidance can be viewed as a limiting form of the discrete guidance in discrete time settings. In addition, the discrete-time guidance has a similar challenge to the rate-based guidance; that is, it requires multiple function evaluations in each sampling step. For solving this problem, Vignac et al. (2023); Schiff et al. (2025) employ the first-order approximation in practice. As pointed out in Remark 1, such approximation-based approaches are inappropriate for discrete data intuitively.
@@ E.3 Comparison between Posterior-Based Guidance and Rate-Based Guidance
Theorem 1 and Theorem 2 present the discrete guidance based on the posterior and the transition rate, respectively. The later one generalizes the discrete guidance of Nisonoff et al. (2025), which requires a stronger condition than the posterior-based guidance in Theorem 1. To better understand the conditions in both theorems, we will analyze the relationship between discrete diffusion and discrete flow models. In the following, we consider two different states $x\neq z$.
Both Guidance Yield the Same Transition Rates in Discrete Diffusion Models. Consider a discrete diffusion model with the corruption transition rate $Q^{q}_{t}(x,z)$. Here, we assume that $q_{1}$ is the density of the data distribution. Following Campbell et al. (2022), the time reversal of $Q^{q}_{t}(x,z)$ is
@@
Since $u_{t}^{q}(z,x|x_{1})\overset{\triangle}{=}Q^{q}_{t}(x,z)\frac{q_{t|1}(z|x_{1})}{q_{t|1}(x|x_{1})}$ is a conditional transition rate that can generate the conditional path $q_{t|1}$ (see Appendix H.1 of Campbell et al., 2024), the discrete diffusion models can be viewed as a special case of discrete flow models. Therefore, the condition $Q^{q}_{t}(x,z)=Q^{p}_{t}(x,z)$ as in Theorem 2 actually implies that we take the same specific conditional transition rate $u_{t}^{p}(z,x|x_{1})=u_{t}^{q}(z,x|x_{1})=Q^{q}_{t}(x,z)\frac{q_{t|1}(z|x_{1})}{q_{t|1}(x|x_{1})}$ in discrete flow models.
They Might Offer Different Marginal Transition Rates for Sampling from Target Distribution. Suppose that we are given conditional probability path $q_{t|1}=p_{t|1}$ and conditional transition rate $u_{t}^{q}(z,x|x_{1})=u_{t}^{p}(z,x|x_{1})$. What we want to know is whether there exists a common corruption transition rate $Q_{t}(x,z)$ such that $u^{q}_{t}(z,x)$ and $u^{p}_{t}(z,x)$ are the reverse transition rates of $Q_{t}(x,z)$ under the data distributions $q_{1}$ and $p_{1}$, respectively. If exists, by the time reversal formula (see, e.g. Sun et al., 2023; Lou et al., 2024), we have
@@
where the third equation we use $\frac{q_{1|t}({\mathbf{x}}_{1}|x)}{p_{1|t}({\mathbf{x}}_{1}|x)}=\frac{q_{1}({\mathbf{x}}_{1})p_{t}(x)}{p_{1}({\mathbf{x}}_{1})q_{t}(x)}$ in equation 15 and the fourth equation we use the marginalization trick in the concrete score matching (Meng et al., 2022; Lou et al., 2024). This completes the proof.